This problem does not tell me how many tickets were for adults and how many for a child. How can I tell this?
solve and check your answer
Tickets for a play at the community theater cost $12 for an adult and $4 for a child. If 130 tickets were sold and the total receipts were $840, how many of each type of ticket were sold?

Answer by stanbon(48557)

Tickets for a play at the community theater cost $12 for an adult and $4 for a child. If 130 tickets were sold and the total receipts were $840, how many of each type of ticket were sold?
------------------
Let # of adult tickets be "a"; Let # of child tickets be "c".
--------------------
Value EQUATION: 12a + 4c = 840
Number EQUATION: a + c = 130
------------
Multiply the 2nd equation by 4 to get:
4a+4c = 520
-----------------
Substract that from the 1st equation:
8a = 320
a = 40 (# of adult tickets sold)
-------------
Substitute to solve for "c":
40 + c = 130
c = 90 (# of child tickets sold)
=====================
Cheers,
Stan H.